# Incoherent Beliefs & Inconsistent Actions In Large Language Models

### v2

@@ Abstract
Real-world tasks and environments exhibit differences from the static datasets that large language models (LLMs) are typically evaluated on. Such tasks can involve sequential interaction, requiring coherent updating of beliefs in light of new evidence, and making appropriate decisions based on those beliefs. Predicting how LLMs will perform in such dynamic environments is important, but can be tricky to determine from measurements in static settings. In this work, we examine two critical components of LLM performance: the ability of LLMs to coherently update their beliefs, and the extent to which the actions they take are consistent with those beliefs. First, we find that LLMs are largely inconsistent in how they update their beliefs; models can exhibit up to a 30% average difference between the directly elicited posterior, and the correct update of their prior. Second, we find that LLMs also often take actions which are inconsistent with the beliefs they hold. On a betting market, for example, LLMs often do not even bet in the same direction as their internally held beliefs over the underlying outcomes. We also find they have moderate self-inconsistency in how they respond to challenges by users to given answers. Finally, we show that the above properties hold even for strong models that obtain high accuracy or that are well-calibrated on the tasks at hand. Our results highlight the difficulties of predicting LLM behavior in complex real-world settings.
@@ 1 Introduction
Figure 1: An example of an LLM betting on the opposite side of its belief.
Large language models (LLMs) have shown rapid improvement in capabilities across a wide range of fields involving real world impact, often with high stakes attached to correctness, such as medical diagnosis, financial decision making, and software engineering. In such scenarios, it is often the case that users only compute metrics – such as accuracy or calibration – on a static and non-interactive test set, but this procedure may give limited insight into how the LLMs will behave in a deployment environment. For example, in a medical setting, measurements can be made on an offline diagnosis dataset; but in a real-world clinical setting, new evidence is sequentially gathered (as new tests are conducted, patient history is taken, patients exhibit changes in their symptoms, etc.), and new actions must be taken based on that evidence. Designing a full simulation of the clinical setting in such cases can be expensive, and may simply not be possible due to privacy and ethical concerns. Yet, it is important to know that LLMs will behave reasonably in such settings.
Two important aspects of ensuring the predictability of LLM behavior in such sequential real-world settings are 1) that they update their prior beliefs coherently when new evidence is introduced and 2) that their actions and recommendations are consistent with their beliefs. In this work, we examine both of these behavioral components. For the former, we measure the extent to which LLMs’ belief updates adhere to Bayes’ rule. We find that in general, LLMs exhibit significant deviation in their updates from Bayes-optimality, implying an internal inconsistency in their world models. Remarkably, we also find that models are such poor updaters of their beliefs that their prior beliefs are often better than their posterior beliefs even though the latter are conditioned on extra information.
For the latter, we perform tests under two designs. First, we elicit the confidences of LLMs on open questions from a prediction market; we then test whether the LLMs actually bet in line with their beliefs on such markets (see Fig.˜1). We find significant deviation in the betting behavior from their elicited beliefs; indeed, in many cases, the LLM does not even bet directionally consistently with its beliefs. Second, we test whether LLMs defend their answers when a user questions them consistently with respect to their confidences (see Fig.˜2). Given an initial answer, the user may cast doubt on that answer; consistent behavior in this setting would entail that LLMs defend their confident answers more frequently than less confident answers. We find that LLMs display moderate self-consistency in this behavior overall, but with significant variability across models and datasets.
Further, we examine the extent to which the above LLM behaviors are correlated to their performance and calibration on the related static task. Our results demonstrate that the action-belief discrepancy is not highly predictable by task performance, and surprisingly in some cases is even negatively correlated with how well calibrated the model is on the task.
Our work provides, to the best of our knowledge, the first comprehensive and multi-faceted study of LLM consistency, especially the coherence of their beliefs and the actions they take with respect to those beliefs, demonstrating the difficulties that may be inherent in evaluating LLMs deployed in real-world interactive settings.
Figure 2: An example of deference-inconsistency in models. Models may stick to answers that they have low confidence in, yet switch for answers with higher confidence.
@@ 2 Related Work
Confidence elicitation and calibration. Extensive recent work has focused on methods for measuring the confidence of LLMs, including logit-analysis (lin2022teachingmodelsexpressuncertainty), sampling-based methods (kuhn2023semanticuncertaintylinguisticinvariances; xiong2024llmsexpressuncertaintyempirical), verbal elicitation (lin2022teachingmodelsexpressuncertainty; xiong2024llmsexpressuncertaintyempirical), and linear probe readouts (azaria2023internalstatellmknows), among others. Further work focuses on methods for improving the calibration of LLM confidences (kadavath2022languagemodelsmostlyknow; kapoor2024largelanguagemodelstaught; cherian2024large; kong2020calibratedlanguagemodelfinetuning). Our work examines LLM consistency behaviors across a variety of confidence elicitation methods; our experimental designs can be extended to any elicitation method. We further find undesirable LLM behaviors whose incidences we show are not strongly correlated to how well-calibrated the LLM is.
Bayesian belief updating. Updating prior beliefs given new evidence in line with Bayesian principles is necessary to behave optimally in a dynamic sequential environment (thompson1933likelihood; ghavamzadeh2015bayesian). Extensive research in human cognition has determined that human reasoning approximately aligns with Bayesian principles, albeit with consistent biases such as base-rate neglect, misweighting of priors, and conservatism in belief updating (Griffiths2006-ov; Griffiths2011-pi; Barbey2007-vk). The Bayesian Brain Hypothesis suggests that human cognition fundamentally operates on Bayesian principles (Knill2004-ir), while Bayesian Theory of Mind models demonstrate how humans reason about others’ beliefs, desires, and social relations (Baker2017). We investigate whether similar principles govern LLM belief updates. Very recently, the concurrent work of imran2025llmbeliefupdatesconsistent probes a similar question.
@@
LLMs as forecasters. Recent work (chang2025llm4ts; tang2024timeseriesforecastingllms) has examined the ability of LLMs to act as time-series forecasters, finding strong predictive performance in both zero-shot and fine-tuned settings. Unlike those works, we are focused not on the accuracy of LLMs as forecasters, but instead probe these models similarly to investigate whether LLMs take actions that correspond with their beliefs about future events.

### v3

# Knowing What You Know Is Not Enough: Large Language Model Confidences Don’t Align With Their Actions

@@ Abstract
Large language models (LLMs) are increasingly deployed in agentic and multi-turn workflows where they are tasked to perform actions of significant consequence. In order to deploy them reliably and manage risky outcomes in these settings, it is helpful to access model uncertainty estimates. However, confidence elicitation methods for LLMs are typically not evaluated directly in agentic settings; instead, they are evaluated on static datasets, such as Q&A benchmarks. In this work we investigate the relationship between confidence estimates elicited in static settings and the behavior of LLMs in interactive settings. We uncover a significant action-belief gap – LLMs frequently take actions that contradict their elicited confidences. In a prediction market setting, we find that models often bet against their own high-confidence predictions; in a tool-use setting, models fail to reliably invoke information-seeking tools when their internal confidence is low; and in a user-challenge setting, models change their answers when they have high confidence in them, whilst sticking to answers they have low confidence in. Crucially, we show that static calibration is an insufficient predictor of consistency in the above dynamic settings, as stronger, better calibrated models are somtimes less consistent than their smaller and weaker open-source counterparts. Our results highlight a critical blind spot in current evaluation methodologies: ensuring that a model knows what it knows does not guarantee that it will act rationally on that knowledge.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: The 3 main experiments of our paper, each showcasing an action-belief gap. Left (Design 1): When asked to bet to maximize a given utility function, LLMs bet inconsistently with – often in completely the opposite direction to – their elicited confidences. Middle (Design 2): In a tool call setting, when provided with a search tool to use to check their answers, LLMs fail to invoke the search despite having low confidence in their answer; and conversely, they sometimes invoke the search despite having high confidence. Right (Design 3): In a user-interaction setting, LLMs stubbornly defend answers which they have low confidence in, but change their minds when they have high confidence in an answer.
Large language models (LLMs) have shown rapid deployment across a wide range of real-world applications with high stakes attached to correctness, such as medical diagnosis, financial decision making, and software engineering (Zhou et al., 2025; Liu et al., 2025a; Chen et al., 2025; Jimenez et al., 2024). Consequently, a large body of work has studied the problem of extracting confidence estimates of LLMs. Such works propose a broad spectrum of confidence elicitation methods, including sampling-based approaches, logit- and likelihood-based measures, and explicit verbalized confidence estimates, among others (Kadavath et al., 2022; Kapoor et al., 2024; Tian et al., 2023; Kuhn et al., 2023).
Despite this diversity of approaches, the evaluation of confidence elicitation methods has remained largely uniform. Methods are typically assessed via measures of calibration – most commonly, expected calibration error (ECE) – computed on fixed datasets such as question-answering benchmarks. However, as LLM-based systems evolve from passive chat interfaces into active agentic systems capable of executing multi-step workflows, taking actions, invoking tools, and responding strategically to user feedback, continued reliance on these metrics is problematic. One issue is the distribution shift inherent to switching from a static to an agentic setting. In the static setting, the context of the model usually consists solely of the question, presented in isolation; whilst in the agentic setting, it will often have the traces of past interactions – with the environment, user, and/or with the model’s own previous decisions; and moreover, such content may be noisy and imperfect. Very recent work has now started to investigate the effect of this distributional shift on LLM behaviors with respect to their static confidence estimates (Duan et al., 2025). However, there is a second issue with utilizing traditionally derived confidence estimates in the agentic setting; in doing so, one is making the implicit assumption that the LLM acts in line with its beliefs.
In this work, we study this second issue directly. We ask: to what extent do LLMs take actions that are aligned with their own confidence estimates? Across a range of settings, we find that their actions are frequently misaligned. We identify a systematic action–belief gap, wherein models take actions that are inconsistent with what would be rational under their elicited confidences.
We demonstrate this phenomenon across three experimental setups. First, in a utility-maximization setting, we elicit LLM confidences about future propositions. Then, the LLMs are asked to place bets given market odds. We find that models do not place bets in line with their elicited beliefs, with a striking level of divergence – models often place bets in the opposite direction of their expressed high-confidence beliefs. Second, in a simple tool-use setting, models are given access to an oracle tool that guarantees a correct answer, yet frequently fail to invoke the tool even when their elicited confidence in their own answer is near zero. Third, in a user-challenge setting, we observe inconsistencies when handling interactive feedback: models sometimes defer to a user’s challenge when their stated confidence is high, while stubbornly defending their answer when their stated confidence is low.
Having observed the action-belief gap consistently across experimental designs, elicitation methods, and model families, we then analyze whether the degree of such inconsistency is correlated with model strength, task capability, or its calibration quality on the dataset/task at hand. Surprisingly, we find inconsistency is not completely explained by any of these. In particular, we observe well-calibrated closed-source models such as Gemini 2.5 Pro sometimes behaving more inconsistently than much smaller and weaker open-source models. We therefore posit that the action-belief gap represents an orthogonal, and hitherto understudied, component of LLM capability measurement.
In summary, the main contributions of our work are:
  1. 1.
    We devise three simple experimental settings, covering utility maximization, tool use, and user interaction, to measure the degree to which LLMs act rationally in line with their estimated confidences.
  2. 2.
    We perform the above experiments on 7 different models, from 5 model families, including open and closed-source models, and with 3 different confidence elicitation methods. We find that in all cases, models act divergently from their estimated confidences – sometimes, significantly so.
  3. 3.
    We perform further analysis on the observed action-belief gap which shows that it is not fully explained by either task performance of the model, nor its calibration on the task. Further, we find that lower ECE confidence elicitation methods do not necessarily result in more consistent behavior.
2 Experimental Setup
The two main desiderata of our experimental designs are:
  1. 1.
    The designs should be analogues of commonly used real-world challenges and use-cases of LLMs.
  2. 2.
    The ‘active’ variants of the static datasets should be simple, introducing only the required minimum of complexity in order to satisfy Point 1 above.
We perform our experiments on three open-source instruction-tuned language models: Llama 3.1 8B Instruct (Grattafiori et al., 2024), Gemma 2 9B IT (Team et al., 2024), and Mistral Small Instruct 2409 (Mistral AI, 2024), as well as four closed-source instruction-tuned language models: GPT-4o, GPT-4o Mini, Gemini 2.5 Pro, and Gemini 2.5 Flash. These models encompass a range of different sizes, as well as differing pretraining and post-training methods.
In order to assess the beliefs of these LLMs, we use three methods common in previous literature: logit analysis, sampling confidence, and verbal elicitation. For further details of these methods, see Appendix A. We use multiple methods for each experiment as there is no consensus on the single best method for obtaining LLM confidences (Kapoor et al., 2024; Liu et al., 2025b).
3 Experimental Design 1: Utility-Maximization
Our first experimental design tests directly whether LLM confidences of a given proposition translate to the actions they take, by asking the models to maximize their utility given a prediction market.
3.1 Warmup: Coin Toss
As a warm-up, and to ensure that models have the capability of performing the task, we begin with a simplified setting of a market over a fair coin toss. An abridged version of the prompt is given below, with the full version given in Appendix B.
There is a market with binary outcomes of Yes or No to the question: “Will a fair coin that is tossed land heads?” The market has a probability of 0.250 for Yes, and a probability of 0.750 for No. You start with a capital of \$100. Assume your utility function is linear, and you are maximizing your utility. State how much of your \$100 capital you are going to bet in the format: ‘My bet is x on y‘ where x is the amount you wish to bet and y is the side of the market you are taking.
We execute the above procedure with market probabilities for heads set at 0.250 or 0.750, and examine both the linear utility and logarithmic utility cases  [^1] These utility functions both permit simple closed-form expressions for the optimal bet amount.. We then assess the models based on the distance from the optimal bet. We also assess their directional consistency with the optimal bet: this simply denotes whether they bet on the same side of the market as the optimal bet. This metric eliminates the potential confounder of models simply being poor at sizing their bets appropriately.
Table 1: Bet distance from optimal, and directional consistency (in parentheses), for linear and log utilities on a fair coin toss. A distance of 0 is optimal, and 200 is furthest possible from optimal. Consistency ‘Y’ indicates betting correctly on the favorable side.
Linear Utility Log Utility
Implied $P(H)$ Implied $P(H)$
Model 0.25 0.75 0.25 0.75
GPT-4o 0 (Y) 0 (Y) 0 (Y) 0 (Y)
GPT-4o mini 0 (Y) 0 (Y) 8 (Y) 21 (Y)
Gemini 2.5 Pro 0 (Y) 0 (Y) 0 (Y) 0 (Y)
Gemini 2.5 Flash 0 (Y) 0 (Y) 0 (Y) 0 (Y)
Mistral 0 (Y) 0 (Y) 17 (Y) 18 (Y)
Llama 50 (Y) 40 (Y) 33 (Y) 33 (Y)
Gemma 200 (N) 50 (Y) 24 (Y) 94 (N)
Results. Our results are shown in Table 1. We see that GPT-4o and both variants of Gemini 2.5 are perfect in executing this task, and models such as GPT-4o mini and Mistral are also perfect in the linear case, and have reasonably close adherence in the logarithmic utility case. All models are also perfectly directionally consistent, except Gemma, which appears to struggle significantly in this design. The results above indicate that nearly all of our tested models are capable of performing the core task reasonably well.
Figure 2: Mean distance from optimal betting for each model when prompted to maximize either linear or log utility, reported for logit and verbal confidence elicitation. Distances are plotted against expected distances for a no betting baseline (dashed, black) and a 50% probability betting baseline (dotted, gray). Most models perform worse than baseline.
3.2 Prediction Market
We now perform the full experiment, on the Metaculus prediction market, using the dataset described in Appendix E. In this setting, the LLM is required to also synthesize its own beliefs over the outcomes. We use questions that opened after the knowledge cutoff date of all models tested. We elicit confidences via both verbal and logit-based approaches (see Appendix A). An example prompt for logit confidence elicitation is as follows:
We then ask the LLMs to bet on the market, with the actual market implied probabilities provided, with the goal of maximizing a given utility function (linear or logarithmic) under a given capital constraint. We stress to the model in the prompt it should bet in line with its beliefs. An example prompt is given in Appendix C.
Results. Our results are shown in Figure 2. As in the warmup coin-toss setting, we report the average distance of the bets to the optimal bet given their elicited confidences. A model that bets exactly in line with its belief should have a distance of 0. For reference, we also compute the average distances that two baselines would obtain – the ‘no betting’ baseline, which always bets $\mathdollar 0$, and the $50\%$ probability’ baseline, which bets in line with a 50/50 belief over the outcomes. We see that for most models, the average betting distance to their own beliefs is higher than both these baselines for both logit and verbal confidences for logarithmic utility.
As with the coin toss experiment, we also examine whether models bet directionally consistently with their beliefs. We see in Figure 3 that models often bet directionally inconsistently with their beliefs; in no scenario do models achieve more than a $79\%$ match rate, and many strong models such as the GPT series exhibit inconsistency a majority of the time. We further verify that the correlation for each model’s betting directions between the linear and logarithmic settings is around $90\%$, implying that models are self-consistent in their actions, but that these actions are not consistent with their elicited confidences.
4 Experimental Design 2: Tool-Use
In our next experiment, we assess whether LLMs appropriately call tools in line with their confidence estimates. Being able to adeptly perform tool calling is a necessity to achieve strong LLM performance in many agentic settings. Despite this surge in interest, existing benchmarks and evaluations of tool use tend to focus on whether a model can successfully generate a syntactically correct function call or whether it improves task performance after invocation, without explicitly tying the invocation decision to the model’s own uncertainty about its output.
In this experimental design, we directly test whether low confidence in a model’s answer correlates with a rational decision to invoke an available external tool. We present the LLM with a fact-based question taken from the ‘no-context’ subset of TriviaQA (Joshi et al., 2017) and obtain its confidence in its given answer, using both verbal and logit elicitation (see Appendix A). An example question is:
What beverage did Pope Clement VIII officially recognize as a Christian drink in an edict issued in 1592?
Then, in a separate interaction, we present the same question, but we additionally append the following to the prompt:
If you are not sure of the answer, instead of providing it, you may use the tool search(“TEXT TO SEARCH”), which will give you reliably correct answers. Use this tool only if you are unsure of your answer.
Full prompts are given in Appendix D. Further details of the dataset construction are provided in Appendix E.
To assess the consistency of tool call use with respect to confidences, we adopt the position that there is no single ‘correct’ level of uncertainty for an LLM to resort to tool search; such a level may differ between models, due to differences in post-training pressures, or due to differing interpretations of the phrase ‘unsure of your answer’. Instead, we note that consistent models should use the tool more frequently when their confidence is low, and conversely, less often when they have high confidence in an answer. We operationalize this idea by measuring the monotonicity of the no-tool-call rate vs confidence, plotted across all questions presented to the model, using Spearman’s rank correlation. A score of +1 indicates perfect consistency, and -1 indicates maximal inconsistency. Further details of the calculation and motivation for this metric, are given in Appendix F.
Results. Our results are shown in Figure 4. Across all models tested, we observe that behavior in this tool-use setting is generally only moderately aligned with the elicited confidences. While the correlations are generally positive, suggesting models are at least directionally reasonable, they remain far from the perfectly consistent score of +1 for most models and elicitation methods. Indeed, some model/elicitation pairs, such as Mistral with verbal elicitation or Llama with logits, have effectively 0 correlation in their tool call invocation rate. This result further supports our findings outlined in Section 3, and underscores our concerns that LLMs may exhibit substantial action-belief inconsistencies, especially in agentic or autonomous settings.
5 Experimental Design 3: User Interaction
LLMs are increasingly used as interactive assistants for skilled human experts in a wide variety of domains. In such interactions, users may challenge or question the model’s responses; a consistent model should defend answers it has high confidence in, while being more willing to revise answers held with lower confidence. Such behavior would mirror human epistemic practices and align with the normative principle that confidence should guide belief revision (Yeung and Summerfield, 2012).
To probe this property, we design an experimental protocol measuring the deference-consistency of LLMs. We first obtain the model answer to a question, then respond to the model with a challenge phrase, such as ‘Your answer to the initial question is incorrect’, and we record the LLM’s answer to the challenge phrase. If the answer is the same, we say the model ‘stuck’; otherwise, it ‘deferred’. Separately, we elicit the confidence of the model via logit extraction and sampling  [^2] We do not perform sampling for closed-source models due to resource constraints. (as described in Appendix A) on its initial answer. Consistent models should defer at the same or higher rates for answers where they are less confident; such behavior would support consistent and reliable user interactions.
Figure 4: Average tool call consistency by model, with logit and verbal elicited confidences. +1 corresponds to perfect consistency, and -1 to total inconsistency.
As in Section 4, we measure deference consistency by calculating the monotonicity of the ‘sticking rate’ vs ‘confidence’ function for each model and confidence elicitation method. Further details of our metric calculation, and motivation for this metric, can be found in Appendix F. We evaluate our models across four diverse datasets: Code Execution, SimpleQA, GPQA, and GSM-Symbolic; see Appendix E for additional details.
Figure 5: Average deference consistency across datasets, with logit- and sampling-based elicited confidences. Sampling elicitation was used only on open-source models. +1 corresponds to perfect consistency, and -1 to total inconsistency.
Table 2: Correlations of consistency metrics versus dataset performance measures. Spearman’s rank correlations are calculated between the task performance/calibration and consistency metrics of all models. +1 indicates perfect correlation i.e. higher performance/calibration correlates with higher consistency.
Experimental Design Correlation with Task Performance Correlation with Calibration
Design 1: Utility Maximization (Logits, Linear Utility) 0.54 -0.46
Design 1: Utility Maximization (Logits, Logarithmic Utility) 0.64 -0.25
Design 1: Utility Maximization (Verbal, Linear Utility) 0.11 0.25
Design 1: Utility Maximization (Verbal, Logarithmic Utility) 0.64 0.11
Design 2: Tool Calling (Logits) 0.79 0.46
Design 2: Tool Calling (Verbal) 0.53 0.69
Design 3: User Deference (Logits) 0.24 0.16
Design 3: User Deference (Sampling) 0.55 0.45
Average 0.51 0.17
Results. We now report on the deference-consistency of LLMs across our datasets. Our results are shown in Figure 5. A score of +1 corresponds to perfect deference-consistency, and -1 is complete inconsistency. More detailed breakdowns of the results are given in Appendix I.
We find that models generally exhibit moderately positive degrees of deference-consistency. However, there are distinct differences between the models. For example, Gemma has similar sampling-based deference-consistency to Llama, but its logit-based deference-consistency score is much higher. We also note that Mistral, despite being a much larger model than both of these, does not clearly outperform the other two. GPT-4o and GPT-4o mini clearly outperform all other models in deference-consistency, while the strong Gemini models perform no better than the open-source models.
6 Analysis
In this section, we perform analyses and ablations of the experimental designs introduced in the preceding sections.
6.1 Is the action-belief consistency of LLMs predictable?
Given that we have observed a disparity between model confidence estimates and their behavior in the three preceding experimental designs, even in strong closed-source models, a natural question to be asked is whether the level of consistency is related to model characteristics. To this end, we analyze the correlation of the consistency of the LLM in each of our experiments to a) the performance on the task and b) calibration.
Our results are summarized in Table 2. Details of the methodology used for measuring consistency, task performance, and calibration are given in Appendix G.
We see moderately positive correlations with task performance in most of our experimental designs, though the correlation remains far from perfect, and in some cases – such as Utility Maximization with verbally elicited confidences – is nearly 0. More strikingly, the correlation of the calibration of the models on the tasks with their action-belief consistencies is weak; the average across all our designs is only slightly above 0. In some designs, such as Utility Maximization, we even see a negative correlation between consistency and calibration, implying that better calibrated models and elicitation methods exhibit a greater tendency to act out of line with their confidence estimates.
6.2 Which is the most consistent elicitation method?
In the previous subsection, we analyzed the correlation of task performance/calibration with consistency across models, and within each design. We can also average over models and designs  [^3] For utility maximization, we convert the betting distance linearly to the [-1, +1] range to align with the consistency scores of the other two designs., to obtain an overall per-elicitation-method consistency value. These values are reported in Table 3, along with the overall average ECE of each method.
Logits have the highest ECE and are therefore the least well calibrated, by a significant margin. Both verbal and sampling perform similarly in terms of overall calibration; however, they have markedly different average consistency scores. Our conclusion from this is that, as with the results of Section 6.1, the overall quality of the confidence elicitation method on static calibration is not a good predictor of consistency. As we discuss in Section 8, we therefore posit that action-belief consistency is a separate, relatively orthogonal component of LLM behavior to static calibration.
Table 3: ECE and Consistency metrics across different methods.
Method Avg. ECE Avg. Consistency
Logits 0.428 0.472
Verbal 0.127 0.361
Sampling 0.117 0.614
6.3 Ablation: Prompting
In this section, we conduct an ablation to test the effect of the prompt on the action-belief gap. We focus on the User Deference experimental design, which permits the most flexibility with regard to prompt design, and test on the open-source models. We modify the experimental design such that one of the following three additional clauses is appended to the system prompt:
  1. Prompt Ablation 1
  2. Prompt Ablation 2
  3. Prompt Ablation 3
    “Answer succinctly, without any extended step by step reasoning.”
PA 1 is designed to explicitly coax the LLM to behave more consistently with regards to its confidences. PA 2 examines whether having the LLM provide a verbal confidence at the end of its initial response elicits better deference-consistency. PA 3 is an ablation to determine the impact of limiting chain-of-thought reasoning, which is the default behavior of the models we tested.
Table 4: Change in deference-consistency ($\Delta$) from prompt ablations, averaged across models and datasets.
Elicitation Method $\boldsymbol{\Delta}$PA1 $\boldsymbol{\Delta}$PA2 $\boldsymbol{\Delta}$PA3
Logit-based 0.056 0.009 0.019
Sampling-based 0.120 0.192 0.097
Our results are reported in Table 4, and in more detail in Appendix H; the entries denote the improvement in the consistency metric from utilizing the new prompt, over using the standard system prompt. We find that PA1 generally improves performance across models, particularly for Llama. PA2 is the most effective overall, achieving nearly a +0.2 improvement in consistency under sampling-based elicitation across models and datasets. We also find, intriguingly, that sampling-based consistencies are generally improved by a significantly greater amount than logit-based consistencies by the addition of our prompt ablations. We speculate this indicates that long-form generation is more conducive to guidance from prompting than single per-token probabilities; confirmation of this hypothesis is left to future work.
@@ 7 Related Work
Confidence elicitation and calibration. Extensive work has focused on methods for measuring the confidence of LLMs, including logit-analysis (Lin et al., 2022), sampling-based methods (Kuhn et al., 2023; Xiong et al., 2024), verbal elicitation (Lin et al., 2022; Xiong et al., 2024), and linear probe readouts (Azaria and Mitchell, 2023), among others. Further work focuses on methods for improving the calibration of LLM confidences (Kadavath et al., 2022; Kapoor et al., 2024; Cherian et al., 2024; Kong et al., 2020). Our work examines a variety of confidence elicitation methods; our experimental designs can be extended to any elicitation method.
@@
LLMs as forecasters. Recent work (Chang et al., 2025; Tang et al., 2024) has examined the ability of LLMs to act as time-series forecasters, finding strong predictive performance in both zero-shot and fine-tuned settings. Our work does not focus on the accuracy of LLMs as forecasters, but instead, the extent to which their forecasts (and actions contingent on those forecasts) correspond to their elicited confidences.